

********************************************************************
UARS WEEKLY STATUS REPORT - 01 APR 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM was turned off all of this week.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 19 March 2005 (Orbit 73994, GMT Day 078) through
25 March 2005 (Orbit 74098, GMT Day 084).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 079/1957 74020 ZEPS/NEPS DAY ONLY TO OFF
SUSIM 079/2142 74021 POWER OFF
SOLSTICE 079/2204 74022 FULLTIME TO DAY ONLY
A Reverse-to-Forward Yaw Around maneuver (number 138) was successfully
executed in Orbit 74035 beginning 080/ 20:02:25Z and completing at
080/ 20:39:35Z on 21 March 2005. All instruments were in their off or
standby mode (except for SOLSTICE in day-only). Due to the SA
position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 55.71 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night
load bus voltage was 26.24 volts and the average depths of discharge
(DODs) on Batteries 2 and 3 were approximately 20.74 and 18.52 percent
respectively. The next (forward-to-reverse) YAW is scheduled for 29
April 2005 (DOY 119).
The beta angle ranged between the angles of 12.4 degrees to 13.5
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 65 degrees on 09 April 2005 (DOY 099).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 7.05 to 6.31 degrees C. Additional
battery performance data for each battery on 18 March 2005 and 21
March 2005 are provided in Appendix D.
The clock error ranged from 10.4 to -10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 09
PEM = 04
SUSIM = 01
SOLSTICE = 01
MLS = 02
HRDI = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Four (4) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to monitor the batteries closely. On
Saturday evening PEM was scheduled to go from day-only to off at the
change of GMT day. The FOD was told to notify the lead engineer prior
to executing this due to the recent VT change. The FOD informed the
lead that from the realtime plots the differential appeared to
climbing on battery 2 during the daytime portion of the orbit. This
is an indication that the battery is being overcharged, so the FOD was
told to leave PEM on until the next day and then it would be
reassessed.
Then on Sunday the FOD informed the lead that differential voltage was
now climbing during end of night, so the PEM turn off was executed.
Unfortunately, the FOD did not execute the second half of the
procedure and PEM was left on fulltime as a result. This caused the
differential on battery 3 to go from approximately -230 to -500 mV in
one orbit. The mistake was finally found on night shift that night by
the FOD and the situation was corrected. The differential then
stabilized after several orbits until a few orbits before the yaw on
the next day. At that time, it began working its way back up to -500
mV and the FOT was poised to command VT-8, but on the two orbits
before the yaw the differential went back down to -300 mV, so we
remained in VT-7 for the yaw.
The yaw was executed successfully, but the two orbits after the yaw
yielded battery 3 differential voltages of -593 and -694 mV
respectively. After this second orbit the MPS was commanded to VT-8.
The spacecraft remained in this configuration over night and was not
put back into VT-7 until approximately 081/2000 GMT, due to a max
battery 2 temperature of 14.76C.
The batteries remained in VT-7 most of the week with only one other
short stint of VT-8 from the 083/0047 event until the 083/0537 event.
Then on Friday evening right after the change of the GMT day 085/0027
the MPS was commanded to VT-6 due to excessive C/D ratios on battery
2. This caused the differential voltage on battery 3 to max out the
telemetry value at -700 mV. The change in philosophy came from the
power branch meeting that morning, in which they stated that the
batteries should be operated at the level best for battery 2 and no
longer should battery 3 differential be chased. The FOT ended the
week in this configuration and will continue to monitor the batteries
closely over the weekend.
In the meantime, the FOT has begun to put together a plan to get the
solar array drive moving again. This would reduce the spacecraft
night duration by approximately 20 minutes and alleviate the high
battery 3 differential voltage seen at end of night.
Also this week, software development successfully installed the new
CMS software on the prime CMS. This gives UARS the capability to
properly schedule SMA events. UARS will begin using this
functionality starting on 28 March 2005. However, given the current
battery issues all nighttime events will be terminated early and no
instruments will be turned on until the battery performance has
stabilized.
MEETINGS
The FOT held several meetings this week with the NASA Mission
Director, UARS project scientist and NASA power branch to discuss the
current battery situation and the future of UARS operations.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 April 2005 (DOY 119). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 25 March 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
74020 079/1957 05-075 PEM PEM ZEPS/NEPS DAY ONLY TO OFF
74021 079/2142 05-076 SUSIM POWER OFF
74022 079/2204 05-077 SOLSTICE FULL TO DAY ONLY
74023 080/0033 05-078 MLS DISABLE PMON ACTIONS FOR HEATERS
74026 080/0503 05-079 PEM ZEPS/NEPS OFF
74033 080/1658 05-080 MLS DISABLED MLS BACKUP HEATERS
74035 080/1911 05-081 OBC DISABLE TMON #8
74036 080/2144 05-082 OBC YAW AROUND MANEUVER SCRIPT
74038 081/0041 05-083 MPS VT LEVEL 8
74039 081/0120 05-084 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
74050 081/2001 05-085 MPS VT LEVEL 7
74059 082/1014 05-086 OBC CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
74068 082/2327 05-087 MPS VT LEVEL 8
74071 083/0537 05-088 MPS VT LEVEL 7
74074 083/0939 05-088 MPS BATSOC
74083 083/2324 05-089 OBC/HRDI NOOP HRDI DAY ONLY
74094 084/1707 05-090 HRDI POWER ON TO DAY ONLY
74095 084/1951 05-092 PEM PEM HTR ON/OFF
74096 084/2107 05-091 PEM POWER ON TO DAY ONLY
74097 084/2248 05-094 PSU/MPS HRDI, PEM OFF & VT-6
74097 084/2248 05-093 HRDI HRDI ERR CLEAR
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 25 March 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 25 March 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
74039 081/ CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
74059 082/ CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 25 March 2005
18 March 2005 - GMT Day 077 Beta = 14.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.42/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 336.0/ 84.0 +10.74/+7.09 +24.0/-12.8 27.0 1.10 18.6
3 -33.6/-140.0 +6.11/+3.87 +21.6/-11.2 27.0 1.07 16.9
25 March 2005 - GMT Day 084 Beta = 13.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -2.11/- 3.05 +00.0/+00.0 06.9 0.00 0.0
2 212.8/ 89.6 +13.49/+12.10 +29.6/-14.8 26.7 1.14 18.6
3 00.0/-593.6 +6.44/+ 5.79 +24.0/-11.2 26.6 1.07 16.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
74038 081/0041 LV7 -> LV8
74050 081/2001 LV8 -> LV7
74068 082/2327 LV7 -> LV8
74071 083/0939 LV8 -> LV7
74097 084/2248 LV7 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 25 March 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-020 PEM 74026 080/0503 N/A ZEPS/NEPS turned off
05-021 MPS 74038 080/2332 N/A BAT 3 DIFFV -694 R-HI
05-022 OBC 74083 084/0041 N/A S/W Stored Cmd Error
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-015 MPS 73895 071/1606 N/A PMON grp 31/PMON act 21 trip
05-016 FOT 73926 073/1348 N/A Data dropout; bad weather@Guam
05-017 FOT 73926 073/1423 38922 Data dropout; IP-NOC linework
05-018 FOT 73930 073/2043 N/A PMON grp 31/PMON act 22 trip
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-012 FOT See List Mar 05 30832 Generic Late Acquisition
05-020 PEM 74026 080/0503 N/A ZEPS/NEPS turned off
05-021 MPS 74038 080/2332 N/A BAT 3 DIFFV -694 R-HI
05-022 OBC 74083 084/0041 N/A S/W Stored Cmd Error